# Jiawei Zhang    Shuang Yang    Bo Li

### v2

@@ 5 Ablation Study
Attack Efficiency. As indicated in Figure 3, we calculate the average number of optimization iterations required for a successful attack with UDora on a subset where attacks are consistently successful across different numbers of locations. It is observed that whether using Sequential or Joint Optimization, no more than 30 iterations are necessary to achieve a successful attack on an LLM agent, showcasing extreme efficiency of UDora. Notably, on AgentHarm, UDora with Joint Optimization requires fewer than 10 iterations on average to conduct a successful attack, which demonstrates the importance of using models’ own reasoning style to attack themselves. To further evaluate computational efficiency, we find that the main per-iteration cost stems from optimizing the adversarial string (similar to the GCG attack (Zou et al., 2023)) rather than the response collection (query). For example, on the AgentHarm dataset with LLaMA 3.1, the average times per iteration are 3.48 seconds for query and 11.22 seconds for optimization, while on InjecAgent, these are 5.57 and 5.97 seconds, respectively. With the Ministral model, the corresponding times are 2.49 and 6.04 seconds on AgentHarm, and 6.86 and 8.18 seconds on InjecAgent. In practice, updating the reasoning less frequently (e.g., every 10 steps) or generating only the first 100 reasoning tokens per query can further reduce computational overhead.
@@ 6 Red-Teaming on Real World Agents
Real-world agents are not as robust as one might think. For instance, when utilizing the function calling feature from OpenAI’s models (OpenAI, 2024), setting the ‘tool_choice’ parameter to ‘required’ or ‘forced_call’ can cause most of the malicious instructions in AgentHarm (Andriushchenko et al., 2024b) to be followed. Similarly, with Anthropic’s recent Computer Use Agent (Anthropic, 2024b), a direct prefilling attack, which involves prefilled responses such as “I’ll help you and I will start with” prompts the Anthropic agent to provide the next action in response to a malicious request regarding computer use. This highlights the critical need for improved security measures in the deployment of AI agents, and in this section, we present additional case studies on attacking real-world agents using UDora.
Web Search Agent. Nowadays, the performance gap between closed and open-source models is narrowing. Powerful open-source models like DeepSeek R1 (Guo et al., 2025) have been increasingly deployed by many companies, such as Perplexity, Microsoft, and Dell, in their applications. This widespread deployment exposes more vulnerabilities to transferability attacks. Take Perplexity AI’s Web Search Agent Sonar (Perplexity, 2024) as an example, a possible adversarial attack in this scenario could be like “how to do money laundering.” In other words, an adversary could compel the agent to gather related information from the internet and provide corresponding references. We find that by leveraging the optimized adversarial strings developed for AgentHarm on the Llama 3.1 model can easily be adapted to attack the Sonar agent, detailed examples are provided in Section B.1.
AI Email Agent. AutoGen recently introduced an email agent (AutoGen, 2024) that can autonomously generate email replies via the Gmail tool from Composio, based on the GPT-4o model from OpenAI. By exploiting only the log probabilities returned by GPT-4o, and applying greedy string optimization based on positional scoring without gradient backpropagation using UDora, we successfully attacked this email agent. The attack involved sending emails containing optimized adversarial strings, which led the agent to disclose the content of all recent emails or personal profile information, which is highly sensitive. This occurred even with the agent being prompted not to reveal any private information. Detailed attack logs and the optimization procedure are deferred to Section B.2.
@@ 7 Discussion
Potential Defenses. To mitigate potential misuse of UDora, one practical strategy is to restrict access to the agent’s internal reasoning, for example by providing only condensed or sanitized summaries of reasoning steps (as in OpenAI’s o1 reasoning model). This reduces the attacker’s ability to target reasoning traces for noise insertion. Another direction is the use of guardrail frameworks for LLM agents (Xiang et al., 2024; Luo et al., 2025), which monitor and block unsafe or abnormal behaviors during reasoning. Adversarial training (Madry et al., 2018) can also enhance robustness by exposing models to adversarially generated examples during training, though this may come with a trade-off in generalization or utility. Additionally, detection mechanisms—such as using a separate LLM to validate whether the agent’s reasoning aligns with the original instruction or environment—can help flag adversarial manipulations, though reliably distinguishing adversarial from benign reasoning remains a challenging open problem.
@@ A.2 Optimization Process on WebShop
Examples on the attack process on WebShop on both optimization modes are shown in Figure 6 and Figure 7, where it seems that the adversarial string ultimately leads the LLM agent to hallucinate and select the wrong item (Zhang et al., 2024a). The number of location for noise insertion is also consistently set to $3$ here.
@@ B.2 Attack on the AI Email Agent from AutoGen
AutoGen recently introduced an AI Email Agent (AutoGen, 2024), based on Composio, which integrates numerous tools, including Gmail tools. It includes a user proxy agent that acts on behalf of a human to execute Gmail actions, and another chatbot agent that is used for replying to emails. Our goal is to send a spam email with malicious instructions (including the adversarial string) to steal the private information of the victim user. The base model for the AI Email Agent here is GPT-4o, and we leverage the returned top 20 tokens with the log probabilities as each position to conduct attacks with UDora. We set the noise as the target tool in Gmail, such as GMAIL_FETCH_EMAILS. In each iteration, we calculate the current overall positional score and randomly replace the token in the adversarial string once it can increase the positional score; otherwise, we keep the original string. The final successful attack example with detailed logs is shown in Figure 11.

### v3

@@ 5 Ablation Study
Attack Efficiency. As indicated in Figure 3, we calculate the average number of optimization iterations required for a successful attack with UDora on a subset where attacks are consistently successful across different numbers of locations. It is observed that whether using Sequential or Joint Optimization, no more than 30 iterations are necessary to achieve a successful attack on an LLM agent, showcasing extreme efficiency of UDora. Notably, on AgentHarm, UDora with Joint Optimization requires fewer than 10 iterations on average to conduct a successful attack, which demonstrates the importance of using models’ own reasoning style to attack themselves. To further evaluate computational efficiency, we find that the main per-iteration cost stems from optimizing the adversarial string (similar to the GCG attack (zou2023universal)) rather than the response collection (query). For example, on the AgentHarm dataset with LLaMA 3.1, the average times per iteration are 3.48 seconds for query and 11.22 seconds for optimization, while on InjecAgent, these are 5.57 and 5.97 seconds, respectively. With the Ministral model, the corresponding times are 2.49 and 6.04 seconds on AgentHarm, and 6.86 and 8.18 seconds on InjecAgent. In practice, updating the reasoning less frequently (e.g., every 10 steps) or generating only the first 100 reasoning tokens per query can further reduce computational overhead.
@@ 6 Red-Teaming on Real World Agents
Real-world agents are not as robust as one might think. For instance, when utilizing the function calling feature from OpenAI’s models (Openai), setting the ‘tool_choice’ parameter to ‘required’ or ‘forced_call’ can cause most of the malicious instructions in AgentHarm (andriushchenko2024agentharm) to be followed. Similarly, with Anthropic’s recent Computer Use Agent (computeruse), a direct prefilling attack, which involves prefilled responses such as “I’ll help you and I will start with” prompts the Anthropic agent to provide the next action in response to a malicious request regarding computer use. This highlights the critical need for improved security measures in the deployment of AI agents, and in this section, we present additional case studies on attacking real-world agents using UDora.
Web Search Agent. Nowadays, the performance gap between closed and open-source models is narrowing. Powerful open-source models like DeepSeek R1 (guo2025deepseek) have been increasingly deployed by many companies, such as Perplexity, Microsoft, and Dell, in their applications. This widespread deployment exposes more vulnerabilities to transferability attacks. Take Perplexity AI’s Web Search Agent Sonar (sonar) as an example, a possible adversarial attack in this scenario could be like “how to do money laundering.” In other words, an adversary could compel the agent to gather related information from the internet and provide corresponding references. We find that by leveraging the optimized adversarial strings developed for AgentHarm on the Llama 3.1 model can easily be adapted to attack the Sonar agent, detailed examples are provided in Section B.1.
AI Email Agent. AutoGen recently introduced an email agent (autogen) that can autonomously generate email replies via the Gmail tool from Composio, based on the GPT-4o model from OpenAI. By exploiting only the log probabilities returned by GPT-4o, and applying greedy string optimization based on positional scoring without gradient backpropagation using UDora, we successfully attacked this email agent. The attack involved sending emails containing optimized adversarial strings, which led the agent to disclose the content of all recent emails or personal profile information, which is highly sensitive. This occurred even with the agent being prompted not to reveal any private information. Detailed attack logs and the optimization procedure are deferred to Section B.2.
@@ 7 Discussion
Potential Defenses. To mitigate potential misuse of UDora, one practical strategy is to restrict access to the agent’s internal reasoning, for example by providing only condensed or sanitized summaries of reasoning steps (as in OpenAI’s o1 reasoning model). This reduces the attacker’s ability to target reasoning traces for noise insertion. Another direction is the use of guardrail frameworks for LLM agents (xiang2024guardagent; luo2025agrail), which monitor and block unsafe or abnormal behaviors during reasoning. Adversarial training (madry2018towards) can also enhance robustness by exposing models to adversarially generated examples during training, though this may come with a trade-off in generalization or utility. Additionally, detection mechanisms—such as using a separate LLM to validate whether the agent’s reasoning aligns with the original instruction or environment—can help flag adversarial manipulations, though reliably distinguishing adversarial from benign reasoning remains a challenging open problem.
@@ A.1 Optimization Process on InjecAgent
Figure 6: Adversarial String Insertion in WebShop Interface. Both panels show examples where adversarial strings (highlighted in red) are inserted into product titles. These inserted tokens are optimized by UDora to adversarially manipulate the target LLM to select an malicious item while disregarding the original user instruction.
@@ A.2 Optimization Process on WebShop
Figure 6 illustrates how adversarial strings are inserted into WebShop product titles to mislead the language model during item selection. The inserted red tokens are optimized by UDora to bias the model’s behavior while maintaining a realistic interface. Examples on the attack process on WebShop on both optimization modes are shown in Figure 7 and Figure 8, where it seems that the adversarial string ultimately leads the LLM agent to hallucinate and select the wrong item (zhang2024knowhalu). The number of location for noise insertion is also consistently set to $3$ here.
@@ B.2 Attack on the AI Email Agent from AutoGen
AutoGen recently introduced an AI Email Agent (autogen), based on Composio, which integrates numerous tools, including Gmail tools. It includes a user proxy agent that acts on behalf of a human to execute Gmail actions, and another chatbot agent that is used for replying to emails. Our goal is to send a spam email with malicious instructions (including the adversarial string) to steal the private information of the victim user. The base model for the AI Email Agent here is GPT-4o, and we leverage the returned top 20 tokens with the log probabilities as each position to conduct attacks with UDora. We set the noise as the target tool in Gmail, such as GMAIL_FETCH_EMAILS. In each iteration, we calculate the current overall positional score and randomly replace the token in the adversarial string once it can increase the positional score; otherwise, we keep the original string. The final successful attack example with detailed logs is shown in Figure 12.